I need help with this equation. 5/6c-8=7 Please explain step by step on how you got it.
Answer by stanbon(75887)

5/6c-8 = 7
-----
Add 8 to both sides of the equation to get:
(5/6)c = 15
------
Multiply both sides by (6/5) to get:
c = 15(6/5)
---
c = 3*6
---
c = 18
===========================
Cheers,
Stan H.